OpenStudy (anonymous):

Write the expression as one logarithm \[log(x^3y^2)-2log(xy^{1/3})-3log(\frac{x}{y})\]

OpenStudy (anonymous):

I got it...

OpenStudy (anonymous):

log(x^2*y^2)-log(x*y^(1/2))^2-log(x/y)^3= log(x^2*y^2)-log(x^2*y)-log(x^3/y^3)= log(y)-log(x^3/y^3)= log(1/(x^3*y^2))